IN PURSUIT OF THE FAR-INFRARED SPECTRUM OF CYANOGEN ISO-THIOCYANATE, NCNCS, UN-DER THE INFLUENCE OF THE ENERGY LEVEL DISLOCATION DUE TO QUANTUM MONODROMY

\underline{MANFRED WINNEWISSER}, BRENDA P. WINNEWISSER, IVAN R. MEDVEDEV, FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus Ohio, 43210-1106, USA; STEPHEN C. ROSS, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, P.O. Box 4400, Fredericton NB E3B 5A3, Canada; and JACEK KOPUT, Department of Chemistry, Adam Mickiewicz University, 60-780 Poznan, Poland.

Quantum Monodromy has a strong impact on the ro-vibrational energy levels of chain molecules whose bending potential energy function has the form of the bottom of a champagne bottle (i.e. with a hump or punt) around the linear configuration. NCNCS is a particularly good example of such a molecule and clearly exhibits a distinctive monodromy-induced dislocation of the energy level pattern at the top of the potential energy hump^{a}. The generalized semi-rigid bender (GSRB) wave functions are used to show that the expectation values of any physical quantity which varies with the large amplitude bending coordinate will also have monodromy-induced dislocations. This includes the electric dipole moment components. High level ab initio calculations not only provided the molecular equilibrium structure of NCNCS, but also the electric dipole moment components μ_{a} and μ_{b} as functions of the large-amplitude bending coordinate. The calculated expectation values of these quantities indicate large ro-vibrational transition moments that will be discussed in pursuit of possible far-infrared bands. To our knowledge there is no NCNCS infrared spectrum reported in the literature.

aB. P. Winnewisser, M. Winnewisser, I. R. Medvedev, F. C. De Lucia, S. C. Ross and J. Koput, Phys. Chem. Chem. Phys., 2010, DOI:10.1039/ B922023B